OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21

OpenStudy (anonymous):

For x-intercept: substitute y=0 y-intercept:substitute x=0

OpenStudy (anonymous):

so what is the answer??

OpenStudy (anonymous):

x-intercept:x+7(0)=-21,x=-21 y-intercept:y=-21/7 ,y=-3

OpenStudy (anonymous):

thanks can i ask a couple more question?

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4028/identify-the-correct-intercept-and-intercept-the-equation ask your Algebra2 homework questions at http://www.mathskey.com/question2answer/algebra-2 and get free math help. all the best

OpenStudy (anonymous):

@_tevonn yes

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (bradely):

step by step answers posted here http://www.mathskey.com/question2answer/4030/choose-equation-below-that-represents-passing-through-point ask your Algebra2 homework questions at http://www.mathskey.com/question2answer/algebra-2 and get free math help. all the best

OpenStudy (anonymous):

]Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2 .

OpenStudy (anonymous):

y+4=1/2(x-2) y=(1/2)x-1-4 y=1/2x-5

OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12

OpenStudy (anonymous):

6x-12=3y y=2x-4 slope:2 y intercept :y=-4